Grade 12th passIIT JEE Entrance Exam

A metal surface is illuminated by light of two different wavelengths 248nm and 310nm.The maximum speeds of the photoelectrons corresponding to these wavelengths are U1 and U2 respectively.If the ratio U1 and U2=2:1 and hc=1240eV NM, the work function of material is nearly what?

Anubhav Srivastava
9 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Here, a metal surface is illuminated by light of two different wavelengths 248 nm and 310 nm.
Ratio u₁/u₂ = 4/1 [∴ u₁²/u₂² = 16/1]
hc = 1240 eV nm
=> According to the formula of kinetic energy:
KEmax = hc/λ - w
=> By substituting the value in above formula, we get
For light of 248 nm wavelengths:  
1/2 mu₁² = hc/μ₁ - w
1/2 mu₁² = 1240/248 - w
1/2 mu₁² = 5ev - w ...(1)
For light of 310 nm wavelengths:  
1/2 mu₂² = hc/μ₂ - w
1/2 mu₂² = 1240/310 - w
1/2 mu₂² = 4ev - w ...(2)
=> Dividing eq(1) by eq(2), we get
 
 
16 = 
16(4-w) = 5 - w
64 - 16w = 5 - w
64 - 5 = 16w - w
59 = 15w
w = 59/15
w = 3.93 eV ≈ 3.9 eV
Thus, the work function of the metal is 3.9 eV.